### v2

# thanks: Equal contributionsthanks: Equal contributions

Abstract
Edge states in topological systems have attracted great interest due to their robustness and linear dispersions. Here a superconducting-proximitized edge interferometer is engineered on a topological insulator Ta2Pd3Te5 with asymmetric edges to realize the interfering Josephson diode effect (JDE), which hosts many advantages, such as the high efficiency as much as 73% at tiny applied magnetic fields with an ultra-low switching power around picowatt, and a giant interfering magnetochiral anisotropy with a maximal coefficient $\gamma$ = $1.2\times 10^{9}$ T-1A-1. As an important element to induce such JDE, the second-order harmonic in the current-phase relation is also experimentally confirmed by half-integer Shapiro steps. This edge interferometer offers a novel and effective method to enhance the overall performance of JDE and magnetochiral anisotropy, and boosts great potential applications for future superconducting quantum devices.
The semiconductor diode is a fundamental component in modern electronics due to the non-reciprocal responses 1. Analogous non-reciprocal charge transport in superconductors - namely superconducting diode effect (SDE) - has great potential for superconducting quantum electronics, since Josephson junctions (JJs) and superconducting quantum interference devices (SQUIDs) have been key components of superconducting quantum devices 2; 3. SDE in the JJs - namely Josephson diode effect (JDE) - and in intrinsic superconductors has been theoretically proposed in various systems with broken time-reversal and inversion symmetries 4; 5; 6; 7; 8; 9; 10; 11; 12. Field-induced and field-free SDE/JDE has been experimentally observed in various superconductors 13; 14; 15; 16; 17; 18; 19; 20, supercurrent interferometers 21; 22; 23; 24, and other systems 25; 26; 27; 28; 29. Interestingly, asymmetric supercurrent interferometers or SQUIDs with a non-sinusoidal current-phase relation (CPR) provide a good platform to realize JDE (Fig. 1a) with ultra-small magnetic fields and ultra-low power consumption 22; 21, both of which are crucial elements for applications at ultra-low temperatures. Moreover, it is quite feasible to promote their efficiency by varying the generic configuration 9, which is compatible with state-of-the-art lithography technology, and integrate them into large-scale superconducting quantum circuits.
Figure 1: Mechanisms for SDE/JDE. $\mathbf{a}$, JDE in the asymmetric SQUID. $\mathbf{b}$, An asymmetric SQUID formed by asymmetric edge states in Ta2Pd3Te5 JJ and its device configuration. The upper and lower edges are the palladium and tantalum-tellurium atomic chain, respectively. The inset shows the photomicrograph of device S2, and the white scale bar corresponds to 1 $\mu$m. $\mathbf{c-f}$, Simulated CPR based on the minimal model. JDE only exists in $\mathbf{d}$ with different supercurrents, a non-zero magnetic field ($\phi/\phi_{0}\neq 0$) and higher harmonics in the CPR ($\alpha_{n}\neq 0$). $\mathbf{g}$, Band structure of monolayer Ta2Pd3Te5. Edge states (blue and red lines) exist near the Fermi level and are marked in the corresponding edges of $\mathbf{b}$. $\mathbf{h}$, $\left|I_{b}\right|-\left|V\right|$ curves for positive and negative current sweep at $B_{z}$ = 8.4 mT. $\mathbf{i,j}$, Alternating switching between the superconducting and normal states at $B_{z}$ = 8.4 mT and 10 mK in device S2.
Essentially, the appearance of JDE is related to the difference between maximum($I_{c+}$) and minimum($I_{c-}$) current in the CPR. The conventional CPR is an odd function of superconducting phase difference $\varphi$ between superconducting leads, entailing a zero supercurrent at $\varphi$ = 0, which shifts [meaning $I(\varphi=0)$neq 0] when both time-reversal and chiral symmetries are broken 30; 21; 31. To achieve JDE, other mechanisms to cause the CPR deviating from standard sinusoidal $I(\varphi)$ = $I_{c}$ $sin(\varphi)$ will also be introduced. The characteristics of CPR, which are usually inferred by the interference pattern of SQUIDs and the Shapiro steps due to ac Josephson effect, as well as the newly-discovered JDE with non-reciprocal critical current, have been widely harnessed to detect the anomalous superconductivity 32; 33, such as proximitized helical/chiral topological edge states. In turn, in a SQUID constructed by two asymmetric edge states of the high-order topological insulator (HOTI), JDE with an unconventional CPR could be easily realized with highly-tunable efficiency 5 and relatively easy fabrication (Fig. 1b). The relatively small number of edge supercurrent channels, in comparison to the case of bulk states transport, may contribute to the small critical current and power consumption 33; 34.
More precisely, the interferometer formed by two asymmetric edge supercurrents (Fig. 1b) can be viewed as a SQUID with two different JJs, and induces JDE with requirements similar to the asymmetric SQUID 9 in Fig. 1a. According to the Souto $et$ $al$. 9, a minimal model with $n$ JJs concatenated in an interferometer array is applied to account for the origin of JDE, and the CPR only takes the first and second harmonic contribution into consideration. The current is written as $I(\varphi,\phi/\phi_{0})=\sum_{1}^{n=2}I_{n}sin(\varphi+2\pi(n-1)\phi/\phi_{0})+\alpha_{n}I_{n}sin(2\varphi+4\pi(n-1)\phi/\phi_{0})$ for two JJs in parallel, where $I_{n}$ ($\alpha_{n}I_{n}$) is the amplitude of the first (second) harmonic content for the $n$th JJ, $\alpha_{n}$ is current coefficient of the higher harmonic, $\phi$ is the magnetic flux and $\phi_{0}=h/2e$ is the magnetic flux quantum. In Fig. 1c-f, the simulations of the CPR using the aforementioned formula serves the purpose of illustrating three key requirements that contribute to the difference between $I_{c+}$ (red arrow) and $I_{c-}$ (pink arrow). First, an external magentic field ($\phi/\phi_{0}\neq 0$) is necessary to break time-reversal symmetry and induce nonzero current at $\varphi$ = 0 in the CPR (Fig. 1d-f). Second, the supercurrent of each edge, including all the harmonic contributions, should be different (see Fig. 1d,e and detailed simulations in supplementary Fig. S2). This condition can be realized by two interfering edge states with different dispersion 5. It can also be met, for example, by two superconducting quantum point contacts with different transmission coefficients 9, which may be linked to various factors such as different edge states, disorder situations, among others. All of these can cause the different supercurrent of edges in our devices, which is actually quite unique in topological materials. Finally, at least one edge channel should be transmissive so that the CPR acquires a higher harmonic ($\alpha_{n}\neq 0$) 9; 35 (see Fig. 1d, f).
Interfering JDE
In this work, highly efficient JDE, residing in the tilting supercurrent pattern, is reported in a Ta2Pd3Te5 edge interferometer at very small $B_{z}$ and is driven by low power consumption to achieve a highly stable rectification effect. It is also accompanied with fractional Shapiro steps, revealing the higher harmonic contributions of the non-negligible transmission channels and predominantly enhanced interfering magnetochiral anisotropy (MCA). The choice of Ta2Pd3Te5 is based on the following reasons: (1) Ta2Pd3Te5 is a van der Waals material with quasi-one-dimensional (1D) chains 36 and can be easily mechanically exfoliated. (2) As a 2D topological insulator, where multiple layers could be viewed as simple stacking of monolayers due to very weak inter-layer coupling 37; 38, it hosts excitonic insulator states 39; 40; 41 and edge states observed by scanning tunnelling microscopy 42; 40 and transport measurements 38. Notably, owing to the anisotropic bonding energy of 1D chains, the edges possess different atomic chains (Pd and Ta-Te atomic chains), resulting in asymmetric dispersion, as shown by the red and blue lines/arrows in Fig. 1b, g. (3) Superconductivity can be easily achieved in this system through doping or high-pressure techniques 43; 44; 45. Therefore, by proximity effect, the induced supercurrent (supplementary Fig. S1) carried by edge states of Ta2Pd3Te5 can be used to realize an asymmetric SQUID with the interfering JDE.
Figure 2: Asymmetric Josephson effect and JDE in Ta2Pd3Te5 JJs. $\mathbf{a}$, SQUID pattern of device S1 at 15 mK. $\mathbf{b}$, Position-dependent supercurrent density distribution. Inset: A sketch of edge-supercurrents-formed JJ. $\mathbf{c}$, $B_{z}$-dependent $I_{c+}$ and $\left|I_{c-}\right|$. $\mathbf{d}$, Non-reciprocal critical current $\Delta I_{c}=I_{c+}-\left|I_{c-}\right|$, Josephson diode efficiency $\eta$ and fitted $\Delta I_{c}$ (red line) as a function of $B_{z}$. $\mathbf{e}$, Temperature-dependent $I_{c+}$, $I_{c-}$, $\Delta I_{c}$ and $\eta$ for device S1 at $B_{z}$ = 6.4 mT. $\mathbf{f}$, SQUID pattern of device S2 at 10 mK. $\mathbf{g}$, Supercurrent density distribution. The inset is a draft of the JJ with three edge supercurrents. $\mathbf{h}$, $B_{z}$ dependence of $I_{c+}$ and $\left|I_{c-}\right|$. $\mathbf{i}$, Oscillating $\Delta I_{c}$, $\eta$ and fitted $\Delta I_{c}$ (red line). $\mathbf{j}$, Temperature-dependent $I_{c+}$, $I_{c-}$, $\Delta I_{c}$ and $\eta$ for device S2 at $B_{z}$ = 8.4 mT.
[…]
The superconducting half-wave rectification is observed at $B_{z}$ = 8.4 mT and 10 mK in Fig. 1i, j. The ‘on’/‘off’ (superconducting/normal) states are switched by alternating current bias ($I_{b}=\pm$400 nA). Many cycles of these alternating measurements are conducted over a measurement time exceeding 1.5 hours, demonstrating the high stability of this rectification device, which is crucial for its potential applications. Intriguingly, the estimated switching power ($I_{b}^{2}R_{N}$) reaches the picowatt level (6.4 pW for device S2 with $R_{N}=40$ $\Omega$, 0.56 pW for device S1 in supplementary Fig. S1), which is four and eight orders of magnitude smaller than the field-free Josephson diode 25 and another bulk superconudcting diode 13, respectively. This power is also close to or even lower than the power of nanowire/nanoflake SDE systems 21; 19; 46; 47 and thin film SDE systems 15; 48; 18. The ultra-low switching power in Ta2Pd3Te5 Josephson diodes may be attributed to fewer edge supercurrent channels, making it a potential candidate for applications. Nonetheless, further technological advancements are required to optimize the efficiency, magnetic field conditions, and power consumption not only in this Josephson diode, but also in other superconducting diodes.
To further study the origin of JDE, we measure the magnetic flux dependence of the JDE in the Ta2Pd3Te5 JJ. In Fig. 2a, $dV/dI$ as a function of $B_{z}$ and $I_{b}$ displays a SQUID pattern for device S1, suggesting the supercurrent interference of two channels in this JJ. The sawtooth-shaped $I_{c}$ pattern shows up, usually indicating the involvement of higher harmonic components due to non-negligible transmission channels 49, consistent with findings from the supplementary section I. The position-dependent supercurrent density distribution (Fig. 2b) extracted from the SQUID pattern 33 and the estimated enclosed area formed by two interference supercurrents (see supplementary section III) illustrate the supercurrent mainly originates from edge states of the JJ, confirming the reported edge state in Ta2Pd3Te5 42; 38; 40. In Fig. 2c, the $B_{z}$-dependent $I_{c+}$ and $\left|I_{c-}\right|$ extracted from Fig. 2a shows the asymmetric Josephson effect. Several features, such as the large deviation of $I_{c}$ from zero at half flux quantum in Fig. 2c and JDE in Fig. 2d, indicate the asymmetric edge supercurrent channels in this JJ, as detailed in methods. Moreover, $\Delta I_{c}=I_{c+}-\left|I_{c-}\right|$ oscillates with $B_{z}$ (orange curve in Fig. 2d), which can be called interfering JDE and is approximately described by the two-JJs model (red line in Fig. 2d, and see the detail in methods). This type of JDE has been theoretically predicted in Weyl semimetals with broken inversion symmetry and asymmetric helical edge states 5. Furthermore, the oscillating Josephson diode efficiency (cyan line in Fig. 2d) shows a maximal efficiency $\eta\approx$ 45 % at 6.4 mT. It is intriguing that $\eta$ can reach 10 % at $B_{z}$ = -0.5 mT. Also, the required field can be significantly reduced when the enclosed area of the SQUID loop is increased, which could be easily realized in the experimental setup 22; 21.
Figure 3: Fractional Shapiro steps under microwave in both Ta2Pd3Te5 devices. $\mathbf{a}$, Differential resistance ($dV/dI$) as a function of voltage characteristics and flux quantum at microwave power = -32 dBm, $f$ = 5 GHz and 30 mK for device S1. The ±$\pm 1/2$th Shapiro steps are marked by black arrows. $\mathbf{b}$, $dV/dI$ versus $I_{b}$ at $B_{z}$ = 5.5 mT (3/2 $\phi_{0}$). Valleys characterized by half-integer steps are observed clearly. $\mathbf{c}$, Flux quantum and voltage dependence of $dV/dI$ at microwave power = 10 dBm, $f$ = 4.64 GHz and 10 mK for device S2. $\mathbf{d}$, Valleys characterized by the $-1/2$th Shapiro step at 6.7 mT (3/2 $\phi_{0}$). $\mathbf{e}$, Microwave power dependence of $dV/dI$ at 4.64 GHz and $B_{z}$ = 6.7 mT in device S2. $\mathbf{f}$, Microwave-power-dependent non-reciprocal critical current of the 0th Shapiro steps, which is extracted from $\mathbf{e}$.
Having demonstrated the JDE in Ta2Pd3Te5 edge interferometer, we now show how to further enhance its efficiency. Interestingly, compared with device S1, the SQUID pattern of device S2 in Fig. 2f reveals two main periods, leading to a more pronounced asymmetry in the pattern (Fig. 2h, i) and an increased efficiency with $\eta_{max}$ = 73 % at 8.4 mT. The enhanced JDE can be explained by three JJs in parallel. The formation of the middle edge state in the JJ may stem from the ladder-like structure induced by imperfect exfoliation (as indicated by the green arrow in the inset of Fig. 2g and supplementary Fig. S3d), which is a common occurrence during the exfoliation process of 2D flakes 50 . These three supercurrent channels form two sets of interference patterns (corresponding red and green dashed loops in the inset of Fig. 2g), leading to the asymmetric SQUID. The $\Delta I_{c}$ can also be approximately fitted by the three-JJs model (red line in Fig. 2i, and see the detail in methods). The optimization of $\eta$ in device S2 (Fig. 2j), compared to device S1 in Fig. 2e, seems to be achieved by increasing the number of JJs in parallel, which is in line with the theoretically predicted method 9.
Here, we examine the main mechanisms to determine which one best fits our data. (1) Magnetism mechanism 20 is initially excluded due to the absence of magnetism in the Ta2Pd3Te5 JJ. (2) Rashba-SOC mechanism is excluded as it typically requires in-plane magnetic fields, and Ising-SOC is also considered implausible because it is usually found in 2D superconductors 51. (3) JDE/SDE caused by finite-momentum Cooper pairing generally also needs in-plane magnetic fields 8; 14. The absence of obvious enhancement of the upper critical field under in-plane magnetic fields or the approximately butterfly interference pattern in our JJs 14 (supplementary Fig. S5) lead to the exclusion of finite-momentum Cooper pairing as a mechanism. (4) Nonlinear capacitance 6; 28 is absent in this system due to the symmetric superconducting electrodes and the lack of obvious hysteresis between switching current and retrapping current (Fig. 1h). (5) Self-inductance is also found to be too small to be considered (see supplementary section VIII). Therefore, the main origin of $B_{z}$-induced JDE in the Ta2Pd3Te5 edge interferometer is attributed to asymmetric edge supercurrents, as detailed in methods.
Fractional Shapiro steps
As discussed above, besides asymmetric edge supercurrents and time-reversal symmetry breaking, the remaining higher harmonic in the CPR is further studied to explain the JDE in our JJs. The Shapiro resonances have been widely used to reveal not only the exotic symmetry of Cooper pairing 32 but also the higher harmonic or non-sinusoidal CPR 9; 52. The Shapiro steps appear in the $I$ $V$ curve at $V=V_{n}\equiv nhf/2e$ ($n$ is an integer) when the JJ is irradiated with the microwave. If the $n$th harmonic contributes to the CPR, $m/n$th Shapiro steps ($m$ is an integer) will be present 9; 52. In Fig. 3a, fractional Shapiro steps emerge at a microwave frequency of $f$ = 5 GHz, indicating the significant transparency of the JJ (supplementary section I). The ±$\pm 1/2$th Shapiro steps at half flux quantum $(n+1/2)\phi_{0}$ are marked by black arrows, which is also observed in other works 53; 24. The characteristic $dV/dI$ valleys formed by half-integer steps can be clearly observed at marked positions in Fig. 3b at $B_{z}$ = 5.5 mT (corresponding to 3/2 $\phi_{0}$). In addition, half-integer Shapiro steps are also observed in device S2 with flux-dependent measurements ($f$ = 4.64 GHz in Fig. 3c, d or 5.02 GHz in supplementary Fig. S7) and power-dependent measurements at $23\phi_{0}/$ 2 (Fig. 3e). These experiments indicate the existence of the second harmonic in the CPR, which is also supported by the rapid suppression observed in temperature-dependent $\Delta I_{c}$ compared with $I_{c}$ in device S1 (see Fig. 2e), because higher harmonics are quickly suppressed when temperature approaches $T_{c}$ 54. The rapid suppression of $\Delta I_{c}$ in device S2 (Fig. 2j) seems less pronounced, possibly because its three-JJs model enhances the diode effect, therefore relatively weakening such suppression. In addition to their applications in Josephson diodes, the JJs with higher harmonics also have potential applications in 0 - $\pi$ qubits, particularly if the first harmonic in the CPR can be effectively eliminated 55; 56; 57.
Furthermore, some features in Fig. 3 require further explanation. First, the disappearance of half-integer Shapiro steps at zero field (Fig. 3a, c) may be attributed to the relatively small contribution of the second harmonic in the CPR and the low microwave frequency 58. In device S1, a higher microwave frequency is utilized at zero field, and a small signal of half-integer steps appears in supplementary Fig. S6d-f. No higher frequency is applied to detect half-integer steps at 0 T, considering the limitations of our microwave generator. Second, the integer Shapiro steps at 5.5 mT (3/2 $\phi_{0}$) appear weaker in Fig. 3b (or supplementary Fig. S6c) compared to those at 0 T in supplementary Fig. S6b. This is due to the application of a magnetic field at half the flux quantum, causing a reduction in the sharpness of the integer steps because of a decrease in the first harmonic of the CPR. Subsequently, the second harmonic or fractional Shapiro steps become visible at half the flux quantum (Fig. 3a), in agreement with theoretical predictions 9; 35 and experimental results 53; 24. Third, in device S2 (Fig. 3c), half-integer steps vanish at ±$\pm\phi_{0}/2$ due to the superposition of multiple SQUID patterns, which is slightly different from the single SQUID pattern observed in device S1. In the case of multiple SQUID patterns, the position of destructive interference for the first harmonic differs from that of a single SQUID pattern, leading to a notable contribution from the first harmonic and the absence of half-integer steps at ±$\pm\phi_{0}/2$ (see supplementary section VI).
Figure 4: Magnetochiral anisotropy of Ta2Pd3Te5 device S2. $\mathbf{a}$, $R_{w}$ versus $B_{z}$ at different temperatures. $\mathbf{b}$, Antisymmetric $R_{2w}$ with respect to $B_{z}$ at various temperatures. $\mathbf{c,d}$, $R_{w}$ and $R_{2w}$ curves at 20 mK. Pairs of antisymmetric peaks ($R_{2w}$) are marked by dashed lines with the same colors. $R_{2w}$ shows the similar behavior to $\Delta I_{c}$ in $\mathbf{d}$. $\mathbf{e}$, MCA coefficient $\gamma$ as a function of $B_{z}$ from 20 mK to 0.45 K. $\gamma$ decreases with increasing $B_{z}$. $\mathbf{f}$, Temperature-dependent $\gamma$. The average $\gamma$ is calculated using pairs of antisymmetric peaks in $\mathbf{d}$. $\mathbf{g}$, Summary of $\gamma^{\prime}$ as a function of $R_{2w}/R_{w}$. Ta2Pd3Te5 interferometer exhibits the huge $\gamma$ at large $R_{2w}/R_{w}$. The open (half-fulled, fulled) scatters represent the JJs (normal conductors, intrinsic superconductors).
The JDE is also studied as a function of microwave power. In Fig. 3f, switching current $I_{c+}^{0th}$ and $I_{c-}^{0th}$ of the 0th Shapiro step are extracted from the Fig. 3e. The quantity $\Delta I_{c}^{0th}=I_{c+}^{0th}-\left|I_{c-}^{0th}\right|$ (orange curve, see the detail in supplementary Fig. S8) displays little variation and gradually decreases with increasing microwave power. This stable observation of the JDE under microwave irradiation represents one of the initial examples of such behavior and is a crucial feature in quantum information applications, where microwaves are commonly utilized 2.
Interfering MCA
As another type of non-reciprocal behavior, MCA is measured to characterize non-reciprocal charge transport 59; 13; 25, with its strength often defined by the coefficient $\gamma=\left|\frac{2R_{2w}}{R_{w}BI}\right|$. MCA typically exhibits antisymmetric behavior in $R_{2w}$ with respect to magnetic fields. In normal conductors, MCA is a small effect ($\gamma$ $<$ 1 T-1A-1), generally due to the small SOC. However, significant enhancement of MCA is observed in noncentrosymmetric superconductors, possibly due to a large difference in the energy scale between the Fermi energy and the superconducting gap 59.
Therefore, it is quite interesting to measure MCA in a supercurrent interferometer. In Fig. 4a, b, the first- ($R_{w}$) and second-harmonic resistances ($R_{2w}$) of device S2 are measured using lock-in techniques. Unlike the linear current dependence of $\left|R_{2w}\right|$ in other systems at small current 60; 61; 62, current-dependent $R_{2w}$ behaviors in our JJs seem a little complex at different fields because of supercurrent interference (supplementary section VII). As a result, an ac current $I_{ac}$ = 500 nA is applied to measure temperature-dependent $R_{2w}$ in Fig. 4 to obtain a large $R_{2w}$ and observe relatively intact features of antisymmetric $R_{2w}$ $B_{z}$ curves. In fact, the oscillating antisymmetric behavior of $R_{2w}$ $B_{z}$ curves in Fig. 4b is different from previous reports 59; 13; 25; 54, where only a pair of antisymmetric peaks exist with respect to the magnetic field. These antisymmetric behaviors can usually provide additional information compared to SDE, including disparities between the Fermi energy and superconducting gap, the Rashba parameter and spin-orbit splitting, among other factors 59; 63. The oscillating $R_{2w}$ (red line in Fig. 4d) is analogous to the $\Delta I_{c}$ of JDE (orange line in Fig. 4d, and similar behavior of device S1 in supplementary Fig. S10-12), indicating a common underlying origin. To our best knowledge, such interfering MCA has not been observed before, and apparently reduces magnetic fields at which MCA appears to a large degree.
In Fig. 4e, an extremely large $\gamma$ = 2.2 ×$\times 10^{9}$ T-1A-1 is obtained at tiny $B_{z}$, suggesting that this is the largest value reported to date 54; 61; 62; 13; 25; 60; 64; 65. With increasing $B_{z}$ and temperature, $\gamma$ gradually decreases, as shown in Fig. 4e, f, respectively. In Fig. 4f, the average $\gamma$ is calculated using pairs of antisymmetric peaks (black and red dashed lines in Fig. 4c, d), and the largest $\gamma_{1}$ and $\gamma_{2}$ are $1.2\times 10^{9}$ and $9.6\times 10^{8}$ T-1A-1, respectively. To compare with the MCA in several typical systems, the relationship between $\gamma^{\prime}$ and $R_{2w}/R_{w}$ is summarized in Fig. 4g, where $\gamma^{\prime}$ ($=\gamma A$, where $A$ is the cross-sectional area of the JJ) is also calculated to characterize the MCA, because $\gamma$ seems to show a positive relation with $R_{2w}/R_{w}$ (supplementary Fig. S9). Although no evident relationship between $\gamma^{\prime}$ and $R_{2w}/R_{w}$ exists in different SDE/JDE systems (Fig. 4g), Ta2Pd3Te5 JJs possess not only the large $R_{2w}/R_{w}$ but also huge $\gamma^{\prime}$. MCA of the device S1 is shown in supplementary Fig. S10-12. The maximum value of $\gamma^{\prime}$ can reach $8.44\times 10^{-5}$ ($1.3\times 10^{-4}$) m2T-1A-1 for device S2 (device S1) with $A$ = $7.03\times 10^{-14}$ ($2.08\times 10^{-14}$) m2, which seems to be the largest value reported so far 64; 65. Furthermore, MCA in interferometers can be greatly enhanced theoretically by increasing the enclosed area of the SQUID loop. Here, the enhancement of $\gamma^{\prime}$ is apparently correlated to the interference of asymmetric edge supercurrents, probably due to the much larger energy difference between the Fermi energy and the mini-gap in Andreev bound states, instead of the bulk superconducting gap 59. In this scenario, the mini-gap is modulated by the flux quantum pinned in the SQUID, which is why $R_{2w}$ oscillates in conjunction with the supercurrent interference. Further investigation of MCA in such types of interference systems is warranted.
Conclusions
The Ta2Pd3Te5 edge interferometer reveals a significant JDE under small out-of-plane magnetic fields, as well as extremely large MCA. Moreover, the presence of asymmetric Josephson critical current alongside fractional Shapiro steps demonstrates the importance of the higher harmonic content in the JDE. It is noted that power consumption requires attention if SDE/JDE is to be applied at low temperatures in the future, and further optimization is necessary not only regarding mechanism, but also in terms of device size. The Josephson diode efficiency can be further enhanced by concatenating interferometer loops. These findings offer a promising and novel approach to exploring JDE with significant diode efficiency, ultra-low switching power at small magnetic fields and stable JDE under microwave irradiation, rendering them potentially valuable for practical applications.
Motheds
Device fabrication. Single crystals Ta2Pd3Te5 were prepared by the self-flux method 36. Ta2Pd3Te5 thin films were obtained through mechanical exfoliating bulk samples onto Si substrates with the 280-nm-thick SiO2 on top, and then coated with PMMA in a glove box at 4000 rpm for 60 s, followed by annealing at 100 C for 120 s. Multi-terminal electrical contacts were patterned after electron-beam exposure and subsequent development. Ti/Al (5 nm/60 nm) electrodes were deposited after Ar etching in order to remove the oxidized layer. After the lift-off step, the device was coated with boron nitride to prevent oxidation. The entire fabrication process took place in a nitrogen atmosphere glove box. The rectangular film, with a long side along the $b$-axis of the crystal, was easily obtained due to its quasi-1D nature 38. The dimensions of device S2 (device S1) are as follows: thickness 37 nm (16 nm), width 1.9 $\mu$m (1.3 $\mu$m), and length 245 nm (400 nm), respectively.
Transport measurements. The electrical transport measurements were carried out in cryostats (Oxford instruments dilution refrigerator). The dc current bias measurements were applied using a Keithley 2612 current source. Both the first- ($R_{w}$) and second-harmonic ($R_{2w}$) resistance were measured through a standard low-frequency (7-11 Hz) lock-in technique (LI5640, NF Corporation). The phase of the first- and second-harmonic signal was set to be 0 and 90, respectively. The current of magnet was applied by a Keithley 2400 current source in order to control the magnetic field accurately. Before the measurement, annealing process with 300 C for $\sim$10 minutes was performed to obtain good contacts in the Ta2Pd3Te5 JJs. $I-V$ curves were obtained though the numerical integration process, and the origin data were all $dV/dI-I_{b}$ curves.
Transport measurements under microwave radiation. The electrical transport measurements under microwave radiation were carried out using PSG-A Series Singal generator (Agilent E8254A). The main microwave frequency between 4.64 and 5.02 GHz was applied due to the large absorption effect of microwave for devices S1 and S2. The dissipation of microwave circuit existed during the measurement of device S2, while another microwave circuit for device S1 was good. So the applied power of microwave was a little large but without inducing obvious thermal effect.
Band structure calculations. The first-principles calculations were carried out based on the density functional theory (DFT) with the projector augmented wave (PAW) method, as implemented in the Vienna ab initio simulation package (VASP) 37. The generalized gradient approximation (GGA) in the form of Pardew-Burke-Ernzerhof (PBE) function was employed for the exchange-correlation potential. The kinetic energy cutoff for plane wave expansion was set to 400 eV, and a 1$\times$9$\times$1 k-mesh was adopted for the Brillouin zone sampling in the self-consistent process. The thicknesses of the vacuum layer along $x$ and $z$ axis were set to > 20 Å.
Simulations of JDE. A minimal model with $n$ JJs in parallel is used to explain the JDE 9, where only the first and second harmonic contributions in the CPR are considered. The supercurrent is approximatively described as $I(\varphi,\phi/\phi_{0})=\sum_{1}^{n}I_{n}sin(\varphi+2\pi\zeta_{n}\phi/\phi_{0})+\alpha_{n}I_{n}sin(2\varphi+4\pi\zeta_{n}\phi/\phi_{0})$, where $\zeta_{n}$ is the modified ratio and $\zeta_{1}$ = 0, $\beta_{2}$ = 1 ($\zeta_{1}$ = 0, $\zeta_{2}$ = 1, $\zeta_{3}$ = 0.23) for the simulation of our two (three) JJs in parallel. $\zeta_{3}$ represents the enclosed-area ratio of two corresponding SQUID patterns in device S2. Preliminary switching currents for different channels are estimated in both devices without considering the second harmonic contribution, in order to simulate CPR and JDE. For device S1, the parameters used for simulating $\Delta I_{c}$ are: $I_{1}$ = 75 nA, $I_{2}$ = 15 nA, $\alpha_{1}$ = 0.3 and $\alpha_{2}$ = 0.5, while for device S2, the parameters are: $I_{1}$ = 330 nA, $I_{2}$ = 120 nA, $I_{3}$ = 158 nA, $\alpha_{1}$ = 0.6, $\alpha_{2}$ = 0.1 and $\alpha_{3}$ = 0.2. The simulated $\phi/\phi_{0}$-dependent $\Delta I_{c}$ can approximatively explain the experimental data in both devices, as shown by red lines in Fig. 2d, i. The slight discrepancy between the simulation and experiment may arise from the omission of higher-order harmonics in the CPR, a reduction in critical current and a potential variation of $\alpha$ with increasing magnetic field, or other contributing factors.
Asymmetric edge supercurrents. We will discuss edge supercurrents, their asymmetry, and the origin of such asymmetry in our cases, using device S1 as an example. Firstly, the SQUID-like pattern observed in Fig. 2a exhibits several interference features of two edge supercurrent channels, as supported by previous reports 42; 38; 40. In this pattern, both the width and amplitude of the center lobe are approximately equal to those of the other lobes. This is obviously contrast to a typical Fraunhofer pattern induced by the bulk transport 33; 34, where the width of the center lobe is twice as large as that of the other lobes, while the amplitude of the center lobe is usually significantly larger than that of the other lobes.
Secondly, there are several features supporting asymmetric edge supercurrents in our devices: (1) Without considering the higher harmonics of CPR, we calculated two edge supercurrent channels ($I_{1}$ and $I_{2}$) from the SQUID pattern in supplementary section III, and in device S1, the ratio between the supercurrents $I_{1}$:$I_{2}$ is approximately 3.5:1 ($I_{1}$:$I_{2}$:$I_{3}$ $\sim$ 6:1:2 for device S2); (2) After considering the second harmonic, from a simple simulation using a two-JJs model in supplementary Fig. S2e, the deviation of $I_{c}$ from zero at half flux quantum alone cannot support asymmetric interference channels. However, the coexistence of this deviation and the JDE can support asymmetric supercurrents, as shown in supplementary Fig. S2d, g, h, and device S1 satisfies this coexistence condition; (3) The position-dependent supercurrent density distribution shows a significantly asymmetric supercurrent distribution. However, it should be noted that such asymmetric supercurrent density distribution can also exist in a symmetric system 33; 34 due to the relatively imperfect experimental data.
Finally, the origin of asymmetric edge supercurrents is further discussed: (1) The DFT calculations in Fig. 1g support the existence of asymmetric edge states, which contribute to asymmetric edge supercurrents and JDE 5; (2) Different transmission coefficients 9 of two edge JJs may also lead to asymmetric edge supercurrents, and can be caused by the different edge disorders, different coupling with the Al contacts, and other factors. Such conditions were not intentionally controlled in our devices.
Discussion of switching power. From a practical perspective, the power consumption of devices should be carefully considered at low temperatures. Although the current cooling power of commercial dilution refrigerators generally reaches the microwatt level at 10 mK 66, a Joule excitation power higher than a picowatt (or nanowatt) may raise the temperature of the device or sensor above 10 mK (or 100 mK) when the mixing chamber is kept at 10 mK 67; 68. The self-heating effect is sometimes noticeable and can also be applied. For example, a microheater with a power output of a few picowatts can generate a temperature gradient as large as several millikelvin per micron, and can be further used to conduct Seebeck or Nernst effect measurements in 2D devices at millikelvin base temperatures 69. Therefore, ultra-low power consumption is essential not only for basic measurements but also for applications in large-scale superconducting quantum circuits, which need ultra-low temperatures to enhance the precision of quantum state measurements.
The edge-induced JDE in Ta2Pd3Te5 may offer distinct advantages in terms of power consumption. The edges of the topological system exhibit a low number of conducting channels compared to bulk states transport. This characteristic can result in a lower critical current and reduced power consumption 33; 34.
Acknowledgments
We thank Jin-Guang Cheng for deep discussions. This work was supported by the National Natural Science Foundation of China (Grant Nos. 92065203, 12174430, U2032204, 11974395 and 12188101), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB33000000 and XDB33030000), the Beijing Nova Program (Grant No. Z211100002121144), the Synergetic Extreme Condition User Facility (SECUF), the Informatization Plan of Chinese Academy of Sciences (CAS-WX2021SF-0102), the Financial Support from Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302500), the China Postdoctoral Science Foundation (Grant No. 2021M703462 and 2021TQ0356) and the Center for Materials Genome. A portion of this work was carried out at the Synergetic Extreme Condition User Facility (SECUF).
Author contributions
J. S. and Y.P.L. conceived and designed the experiment. Y.P.L. and Y.H. fabricated devices with the help of A.Q.W., X.C.G., X.F.S. and Z.K.S.. Y.P.L. and Y.H. performed the transport measurements, supervised by Z.W.D., Z.Z.L., T.Q., G.T.L, F.M.Q., Z.A.X., L.L. and J. S.. D.Y.Y. and Y.G.S. synthesized bulk Ta2Pd3Te5 crystals. H.H.S. and Z.J.W. calculated the band structure. J.K. and J.P.H. provided some supports on theoretical modeling. Y.P.L. and J.S. wrote the manuscript, and all authors contributed to the discussion of results and improvement of the manuscript.
Competing interests: The authors declare no competing interests.